A 50 x 100 mm timber is used as a column with fixed ends. Determine the minimum length required at which Euler’s formula can be used if E= 10 GPa and critical stress up to proportional limit is 30 MPa. What safe load can be carried with a FS of 2 if the length is 2.5 m


1
Expert's answer
2021-01-11T11:33:33-0500

The Euler buckling load is


"P_{cr}=\frac{\pi^2EI}{(KL)^2}\to \sigma_{cr}A=\frac{\pi^2EI}{(KL)^2}"


"L=\sqrt{\frac{\pi^2EI}{K^2\sigma_{cr}A}}"


"I=\frac{100\cdot50^3}{12}=1.04167\cdot10^6(mm^4)=1.04167\cdot10^{-6}(m^4)"


"A=100\cdot50=5000(mm^2)=5\cdot10^{-3}(m^2)"


"L=\sqrt{\frac{\pi^2EI}{K^2\sigma_{cr}A}}=\sqrt{\frac{3.14^2\cdot 10\cdot 10^9\cdot 1.04167\cdot10^{-6}}{0.5^2\cdot 30\cdot10^6\cdot 5\cdot10^{-3}}}=1.65(m)" . Answer


"\sigma_{cr}=\frac{\pi^2EI}{A(KL)^2}=\frac{3.14^2\cdot 10\cdot 10^9\cdot 1.04167\cdot10^{-6}}{5\cdot10^{-3}(0.5\cdot2.5)^2}=13.1(MPa)"


"P_{cr}=\sigma_{cr}A=65500(N)=65.5(kN)"


safe load "65.5\/2=32.75(kN)" . Answer